OpenStudy (anonymous):

I DONT HAVE MUCH TIME TO GET THIS DONE, AND WOULD REALLY APPRECIATE SOME HELP Match the reasons to the corresponding statement to complete the proof. Write the letter of the reason in the box. Here are the statments: Statements Reasons 1) 5(x + 3) + 9 = 59 1) 2) 5x + 15 + 9 = 59 2) 3) 5x + 24 = 59 3) 4) 5x + 0 = 35 4) 5) 5x = 35 5) 6) x = 7 6) Please give the reason for each statment, the choices are below to choose from:) thanks so mu

OpenStudy (anonymous):

A. Reflexive Property B. Symmetric Property C. Transitive Property D. Addition Property of Equality E. Subtraction Property of Equality F. Multiplication Property of Equality G. Division Property of Equality H. Substitution Property I. Additive Identity J. Multiplicative Identity K. Commutative Property L. Associative Property M. Distributive Property N. Given

OpenStudy (amistre64):

what have you been able to fill in so far?

OpenStudy (anonymous):

not really anything, becasue im trying to work on 2 other assignments at the same time, I gota catch up because the semester is almost over for me

OpenStudy (anonymous):

ive been needing to get this done for awhile. and its due in an hour

OpenStudy (sirm3d):

the first one is given.

OpenStudy (amistre64):

you should at least be able to define the first part ....

OpenStudy (anonymous):

im not sure what u mean?

OpenStudy (amistre64):

well, the reason for the first one is simply: GIVEN , since it is what you have been given to work thru

OpenStudy (amistre64):

have you even read thru the options list?

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

if you were to ignore the options list; and just describe what was done in each step in your own words; how would you go about it?

OpenStudy (anonymous):

the second one is a multiplication one i know that lol, but just dont know which multiplication one

OpenStudy (amistre64):

a(b+c) = ab+ac ; we distribute the a thru the paranthesis to get rid of the paranthesis

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

now, how did they get from step 2 to step 3?

OpenStudy (amistre64):

1) 5(x + 3) + 9 = 59 given 2) 5x + 15 + 9 = 59 dist prop 3) 5x + 24 = 59 ???

OpenStudy (amistre64):

forget the option list, use your own words

OpenStudy (anonymous):

I cant use my own words, I have to chose from above

OpenStudy (amistre64):

you have to have an end result from the options list; but as an inbetween step its best to simply make sure you know whats going one from step to step first

OpenStudy (anonymous):

i am thinking that 3 is an adding one?? Im not quite sure.